OpenStudy (anonymous):

How do I do these kinds of problems? given the following functions find f(g(9)) f(x)=1/x-10 and g(x)=1/x

OpenStudy (ranga):

Find g(9) first. Then put that as x in f(x)=1/x-10 and evaluate f(x).

OpenStudy (ranga):

g(x) = 1 / x g(9) = ?

OpenStudy (anonymous):

g(9)=9

OpenStudy (anonymous):

=1/9

OpenStudy (ranga):

Yes. g(9) = 1/9 f(g(9)) = f(1/9) If f(x) = 1/x-10 f(1/9) = ?

OpenStudy (anonymous):

my answer is -89/9

OpenStudy (ranga):

no. x should be replaced by 1/9 not 9.

OpenStudy (ranga):

If a is divide by a fraction b/c it is the same as flipping the fraction and multiplying. a divided by b/c = a * c / b = ac/b

OpenStudy (anonymous):

-9/89:?

OpenStudy (ranga):

f(x) = 1/x-10 Put x = 1/9 in 1/x-10 = 1 / (1/9) - 10 = 1 * 9/1 - 10 = 9 - 10 = -1 f(g(9)) = -1

OpenStudy (anonymous):

that's not one of my answers . Do I have to do anything next?

OpenStudy (ranga):

you probably forgot to put parenthesis in your problem. Did f(x) look likes this:\[f(x) = \frac{ 1 }{ x - 10 }\]or like this:\[f(x) = \frac{ 1 }{ x } - 10\].When writing if you don't use a parenthesis it will be taken as the latter which is what I did.

OpenStudy (anonymous):

the first one

OpenStudy (ranga):

yeah. Then you should write it like this: f(x) = 1/(x-10) The parenthesis tell us that the whole thing is in the denominator. Anyway, from the start the problem looks like this: g(x)=1/x g(9) = 1/9 f(g(9)) = f(1/9) f(x) = 1/(x-10) f(1/9) = 1 / (1/9 - 10) To get rid of the 9 as a fraction, multiply top and bottom by 9 = 9 * 1 / (9 * 1/9 - 9 * 10) = 9 / (1 - 90) = 9/(-89) = - 9/89

OpenStudy (anonymous):

ty! very much!